OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 8 feet and 6 feet and the height is 10 feet?

jhonyy9 (jhonyy9):

so will be the sum of 2*8*10 and 2*6*10 =160+120 =280 sqft